If F1 = 850 N , F2 = 640 N , determine the magnitude of the resultant force.

Determine the direction of the resultant force measured counterclockwise from the positive x axis.
Answer #1

net horizontal force

Fx = 640 cos 30 = 554. 256 N

net vertical force

Fy = 850 - 640 sin 30 = 530 N

net magnitude of force

F^2 = Fx^2 + Fy^2

F = 529.477 N

======

b)

direction

x = arctan ( Fy/Fx)

x = 43.72 above +x axis

======

Comment before rate in case any doubt, will reply for sure.. goodluck